This bird has a mass of 0.8 kg. Calculate the weight of the bird acting downward as it flies
1
Expert's answer
2020-03-23T10:49:38-0400

P=m(g-a)

Since the bird is acting downward as it flies, it's acceleration is equal to g, so a=g, then P=m(g-g)=0